40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Entry  Wed Jul 7 13:05:07 2021, Paco, DailyProgress, 1418 nm AUX ECDL, New aspheric flexures 
    Reply  Fri Jul 9 11:40:42 2021, Paco, DailyProgress, 1418 nm AUX ECDL, New aspheric flexures IMG_1773.jpeg
       Reply  Wed Jul 14 12:21:08 2021, Paco, DailyProgress, 1418 nm AUX ECDL, ECDL lases... and MZ locked mid-fringe selfhomodyne_1418.jpg
          Reply  Wed Aug 4 11:36:30 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL with 2um AOM tests rf_setup.jpgdiffraction_levels.png
             Reply  Tue Aug 10 09:51:44 2021, Paco, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz rf_setup.jpg
                Reply  Thu Aug 12 11:49:59 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz zeroth_order_contour.pdffirst_order_contour.pdf
                   Reply  Tue Aug 17 11:09:29 2021, rana, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz 
                   Reply  Wed Sep 1 13:12:02 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz 
Message ID: 1915     Entry time: Fri Jul 9 11:40:42 2021     In reply to: 1914     Reply to this: 1916
Author: Paco 
Type: DailyProgress 
Category: 1418 nm AUX ECDL 
Subject: New aspheric flexures 

[Radhika, Paco]

We tweaked the flexure alignment until we had a nominally collimated beam (~2 mW @ 250 mA of diode current) through the output aperture in the ECDL housing. We noted that the collimated beam is off-centered on that circular aperture along the horizontal (yaw) angle. After this, Radhika installed the ECDL grating and we hooked up the fiber output onto a InGaAs PD to monitor the power output. We tweaked the alignment of the grating (mostly yaw) to try and see a change in the power output to indicate optical gain in the diode, but saw no changes. We observed a change in the PD photocurrent as a function of the diode current in the absence of the grating (no optical feedback) which is indicative of ASE. We measured this level to be ~ 140 mV at 200 mA of current; with no observed threshold. In conclusion, we still need to refine our grating alignment to provide gain on the diode and observe lasing at the nominal 1450 nm wavelength.

Attachment 1: IMG_1773.jpeg  1.847 MB  Uploaded Fri Jul 9 23:34:05 2021  | Hide | Hide all
IMG_1773.jpeg
ELOG V3.1.3-